OpenStudy (anonymous):

\int \frac{1}{x^{-8}}\, dx

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{x^{-8}} dx\]

OpenStudy (anonymous):

thanks...

myininaya (myininaya):

\[\int\limits_{}^{}x^{8} dx\]

OpenStudy (saifoo.khan):

\[\int\limits \frac{1}{x^{-8}}\] \[\int\limits x^8dx\] \[\Large \frac{x^9}{9}+C\]

myininaya (myininaya):

\[\int\limits_{}^{}x ^n dx =\frac{x^{n+1}}{n+1}+C , n \neq -1\]

OpenStudy (anonymous):

thanks